OpenStudy (mrs.jmr):

A color printer prints 23 pages in 8 minutes. How many pages does it print per minute?

OpenStudy (eliesaab):

23/8=2.875

OpenStudy (eliesaab):

So it cannot print 3 pages per minutes.

OpenStudy (eliesaab):

How many can it print per minute?

OpenStudy (hockeygal11):

Since the printer can print 23 pages in 8 minutes, you must make a ratio to figure out how many it can print in one minute. Set this up like so: 23/8 = x/1 No, you can cross multiply and get this: 8x=23...x=2.875...which means it can only print 2 per minute, since you can't print .875 of a paper.